# Superpixel Graph Contrastive Clustering with Semantic-Invariant Augmentations for Hyperspectral Images

### v1

@@ Abstract
Hyperspectral images (HSI) clustering is an important but challenging task. The state-of-the-art (SOTA) methods usually rely on superpixels, however, they do not fully utilize the spatial and spectral information in HSI 3-D structure, and their optimization targets are not clustering-oriented. In this work, we first use 3-D and 2-D hybrid convolutional neural networks to extract the high-order spatial and spectral features of HSI through pre-training, and then design a superpixel graph contrastive clustering (SPGCC) model to learn discriminative superpixel representations. Reasonable augmented views are crucial for contrastive clustering, and conventional contrastive learning may hurt the cluster structure since different samples are pushed away in the embedding space even if they belong to the same class. In SPGCC, we design two semantic-invariant data augmentations for HSI superpixels: pixel sampling augmentation and model weight augmentation. Then sample-level alignment and clustering-center-level contrast are performed for better intra-class similarity and inter-class dissimilarity of superpixel embeddings. We perform clustering and network optimization alternatively. Experimental results on several HSI datasets verify the advantages of the proposed method, e.g., on India Pines, our model improves the clustering accuracy from 58.79% to 67.59% compared to the SOTA method.
Index Terms:
Hyperspectral image, superpixel, graph learning, contrastive learning, clustering
@@ I Introduction
Hyperspectral remote sensing techniques employ sensors to collect the reflectance of land-cover materials in hundreds of narrow and contiguous spectral bands, generating hyperspectral images (HSI), which have been widely applied in various fields, such as agricultural monitoring[1], mineral exploration[2] and urban planning[3]. Grouping HSI into several classes is essential in many applications[4, 5, 6, 7]. However, due to the high cost of manual annotating, obtaining pixel-level labels is difficult.
Unsupervised clustering methods[8, 9, 10, 11] are gaining increasing attention. These methods aim to mine relationships between pixels in the absence of labels and assign similar pixels to the same cluster. For example, center-based clustering methods, e.g., K-means[12] and fuzzy c-means[13], assume that instances have a spherical cluster structure, which may be not satisfied in the raw feature space of HSI. Subspace-based clustering methods, e.g., sparse subspace clustering[14], low-rank subspace clustering[15], and dense subspace clustering[16], divide instances into different low-dimensional subspaces, which are susceptible to redundant spectral features. Additionally, considering the large quantity of HSI pixels, subspace clustering methods may lead to unbearable computational costs because they need to calculate the self-represented coefficient matrix with time complexity of $\mathcal{O}(n^{3})$, where $n$ is the number of pixels. Therefore, some competitive subspace-based methods are limited to experimenting on a subset of the complete HSI[17, 18].
@@
To overcome the aforementioned drawbacks, we propose a superpixel-level clustering method for HSI with pixel-level pre-training, named superpixel graph contrastive clustering (SPGCC). First, we exploit the 3-D structure of HSI by dividing the original HSI into 3-D cubes. Then, 3-D and 2-D hybrid convolutions are performed on pixel cubes to extract the high-order spatial and spectral features of HSI simultaneously which can be preserved in superpixel-level representations. Next, we aggregate the topological information of superpixels from neighborhoods by GCN. To learn the discriminative and clustering-friendly superpixel representations, we utilize contrastive learning (CL), which brings samples close to their augmented views in the embedding space while pushing different samples’ representations far apart[28, 29, 30]. Proper data augmentations are crucial for CL, as augmented views and original samples should share the same semantics. To this end, we design two effective data augmentations for superpixels, i.e., pixel sampling augmentation and model weight augmentation, to obtain semantic-invariant augmented views. Additionally, traditional CL-based clustering methods treat different samples as negative pairs and push their representations to be far apart from each other[31], even for samples in the same class, which is harmful to clustering accuracy. Differently, we perform sample-level alignment and clustering-center-level contrast on superpixels to obtain better intra-class similarity and inter-class dissimilarity. To mitigate the impact of outliers, we select samples with high confidence to recompute clustering centers. In the whole framework, we alternatively conduct clustering and network optimization, using clustering results to guide the optimization. The proposed SPGCC has an improvement of 8.8%, 2.4%, and 3.2% over the best compared methods on three different HSI datasets.
@@
    We perform pixel-level pre-training to extract high-order spatial and spectral information of HSI, where the pre-training network is completely separated from the subsequent clustering network and can be applied to large-scale HSI.
@@
    We propose a novel superpixel graph contrastive clustering (SPGCC) method and design two semantic-invariant data augmentations for superpixels, i.e., model weight augmentation and pixel sampling augmentation, which can obtain reliable positive samples and improve contrastive learning.
@@
    SPGCC utilizes a more reasonable optimization target compared with traditional contrastive clustering methods and recomputes high-confidence clustering centers considering the outliers. Experimental results on three HSI benchmark datasets demonstrate the effectiveness of our proposed method.
The rest of this paper is organized as follows. Section II briefly reviews preliminary about HSI pixel convolution and graph contrastive learning. Section III presents the proposed SPGCC algorithm in detail, including pixel-level pre-training, graph construction, semantic-invariant superpixel augmentations, and graph contrastive clustering. In Section IV, we provide experimental results and analyses on several HSI datasets. Finally, we summarize our work in Section V.
@@ II-A HSI Pixel Convolution
where $d_{l-1}$ is the number of channels in the $(l-1)$-th layer, $\omega_{i,j,\tau}\in\mathbb{R}^{(2\delta+1)\times(2\gamma+1)\times({2\eta+1})}$ is the kernel’s weight parameter in the $j$-th channel of the $i$-th layer related to the $\tau$-th channel of the last layer, superscripts $(\sigma,\rho,\lambda)$ are indices along corresponding dimensions, $b_{i,j}$ is the bias, and $\phi$ is the activation function.
@@ II-B Graph Contrastive Learning
Several works have applied contrastive learning for graph data due to its remarkable representation ability[37, 38, 39, 40, 41]. Here, we focus on node-level graph contrastive learning (GCL) and divide it into three parts: graph data augmentation, graph encoder network, and contrastive loss function. A graph is denoted as $g=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ is the set of $N$ nodes and $\mathcal{E}$ is the set of $M$ edges. The graph structure is represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ and node attributes are denoted as $\mathbf{X}\in\mathbb{R}^{N\times d}$. We use $\Gamma$ to denote a set of different graph data augmentations and then augmented views are defined as:
@@
where $sim(\cdot)$ is the similarity function and $\tau$ is temperature scaling parameter.
@@ III Proposed Method
Original HSI is 3-D structured and can be denoted as $\mathcal{S}\in\mathbb{R}^{H\times{W}\times{C}}$ with resolution of $H\times{W}$ and $C$ spectral bands. The target is to assign the total $N$ pixels into $K$ classes where $N=H\times{W}$. Our method consists of pixel-level pre-training and superpixel-level contrastive clustering. Since clustering is performed for superpixels, pixels’ labels are inherited from the belonged superpixels. The overall architecture of the proposed method is shown in Fig. 1. First, we employ a variational autoencoder composed of hybrid 3-D and 2-D CNNs for pixel-level pre-training, extracting high-order spatial and spectral features from HSI pixels. Subsequently, superpixel segmentation is performed and the graph structure of superpixels is constructed. In superpixel graph contrastive clustering, we propose two semantical-invariant data augmentations for superpixels and employ sample-level alignment together with clustering-center-level contrast as the target of network optimization. Clustering results are obtained by K-means and then high-confidence clustering centers are recomputed. We conduct clustering and network optimization alternatively.
@@
Encoder $Decoder$
Layer Kernel Output $Layer$ Kernel Output
Input - [-1,1,30,27,27] $Input$ [-1,128]
Conv3D [8,7,3,3] [-1,8,24,25,25] $FC$ [128,256] [-1,256]
BN3D - - $FC$ [256,23104] [-1,23104]
Conv3D [16,5,3,3] [-1,16,20,23,23] $Reshape$ [-1,64,19,19]
@@ III-C Superpixel Semantic-Invariant Data Augmentations
The second superpixel augmentation, model weight augmentation (MWA), is enlightened by the multi-head idea[45, 46]. We design dual branches in the last layer of the GCN encoder to obtain augmented views with semantic consistency. Specifically, the $L$-th layer contains two branches with unshared parameter matrices $\mathbf{W}_{(L)}^{1}$ and $\mathbf{W}_{(L)}^{2}$ which have the same architecture. Due to the different random initialization, two branches’ outputs $\mathbf{H}_{(L)}^{1}$ and $\mathbf{H}_{(L)}^{2}$ represent two different views of the input superpixel but with the same semantic, i.e., a further aggregation result based on the $(L-1)$-th layer.
Combining PSA with MWA, we can obtain different augmented views of the superpixel. Specifically, representations of superpixels $\mathbf{X}^{sp}$ and the corresponding sampled pixels $\mathbf{X}^{p}_{sampled}$ are the inputs of the graph encoder. In the first $L-1$ layers, embeddings of superpixels and sampled pixels are calculated by (11), denoted as $\mathbf{H}_{(L-1)}^{sp}$ and $\mathbf{H}_{(L-1)}^{p}$ respectively. In the last layer, $\mathbf{H}_{(L-1)}^{sp}$ and $\mathbf{H}_{(L-1)}^{p}$ are inputted into both two branches, thus the outputs include four embeddings: $\mathbf{Z}^{sp_{1}},\mathbf{Z}^{sp_{2}},\mathbf{Z}^{p_{1}},\mathbf{Z}^{p_{2}}$ where $\mathbf{Z}^{sp_{1}}$ and $\mathbf{Z}^{p_{1}}$ are the outputs of the first branch for superpixels and sampled pixels, while $\mathbf{Z}^{sp_{2}}$ and $\mathbf{Z}^{p_{2}}$ are of the second. Subsequently, we normalize each row of these embeddings with $l_{2}$-norm.
@@
0:  HSI $\mathcal{S}$; Superpixel segmentation $\mathbf{T}$; Max iteration $I$.
@@ III-D Superpixel Contrastive Clustering
Traditional contrastive clustering methods suffer from low intra-class consistency caused by wrong negative pairs[47]. To overcome this deficiency, we utilize sample-level alignment (SLA) and clustering-center-level contrast (CLC) as optimization objectives. Briefly, at the sample level, we do not treat different superpixels as negative pairs since they may belong to the same class. Superpixel embeddings are only encouraged to be aligned with their semantic-invariant augmented views. At the clustering center level, for a certain clustering center, it is natural to set the other $K-1$ centers as its negative views and perform contrast. Combining SLA with CLC, we hope to get superpixel embeddings with better intra-class similarity and inter-class dissimilarity.
@@
$\mathcal{L}_{CLC}=\frac{1}{K}\sum_{k=1}^{K}-log\frac{exp(\mathbf{c}_{k}^{1}\cdot\mathbf{c}_{k}^{2}/\tau)}{\sum_{j=1}^{K}exp(\mathbf{c}_{k}^{1}\cdot\mathbf{c}_{j}^{2}/\tau)},$ (19)
where $\tau$ is the temperature scaling parameter.
@@ IV-A Datasets
Dataset
Size
#Band #Class #Labeled pixels
India Pines
(145, 145)
200 16 10249
Salinas
(512, 217)
204 16 54129
Pavia University
(610, 340)
103 9 42776
@@ IV-B Experiment Setup
[…]
@@
Dataset
$M$
$\eta$ $\alpha$ $\lambda$
India Pines
1100
[…]
Salinas
2700
[…]
Pavia University
2200
[…]
2) Compared Methods: To comprehensively evaluate the performance and applicability of the proposed SPGCC, we introduce ten baselines for comparison, including K-means[12], fuzzy c-means (FCM)[13], sparse subspace clustering (SSC)[14], scalable sparse subspace clustering (SSSC)[48], selective sampling-based scalable sparse subspace clustering (S5C)[49], superpixelwise PCA approach (SuperPCA)[50], graph convolutional optimal transport (GCOT)[51], neighborhood contrastive subspace clustering (NCSC)[9], superpixel-level global and local similarity graph-based clustering (SGLSC)[8] and dual graph autoencoder (DGAE)[10]. Specifically, for K-means and FCM, we perform clustering at the pixel level. For SSC, SSSC, S5C, clustering is performed at the superpixel level to overcome the high memory cost of the subspace clustering. For other methods, we follow the original settings proposed in their papers. A brief summarization of the evaluated methods can be found in Table IV.
@@
Method
Superpixel
Subspace clustering Graph learning Deep learning
@@
SSC
$\checkmark$
$\checkmark$
SuperPCA
$\checkmark$
SSSC
$\checkmark$
$\checkmark$
S5C
$\checkmark$
$\checkmark$
@@
SGLSC
$\checkmark$
$\checkmark$ $\checkmark$
DGAE
$\checkmark$
$\checkmark$ $\checkmark$
NCSC
$\checkmark$
$\checkmark$ $\checkmark$
SPGCC
$\checkmark$
$\checkmark$ $\checkmark$
@@
3) Clustering Metrics: To quantitatively evaluate the clustering performance, nine popular metrics are utilized: overall accuracy (OA), average accuracy (AA), kappa coefficient (Kappa), normalized mutual information (NMI), adjusted Rand index (ARI), F1-score (F1), Precision, Recall and Purity. Among these metrics, OA, AA, NMI, F1, Precision, Recall, and Purity range in [0,1] while Kappa and ARI range in [-1, 1]. For each metric, a higher score indicates a better clustering result. Similar to the common process of clustering evaluation, the Hungarian algorithm[52] is adopted to map the predicted labels according to the ground truth before calculating metrics. In addition, we log the time cost of each method.
@@
TABLE V: Clustering Performance on HSI Datasets, India Pines, Salinas and Pavia University are Abbreviated to IP, SA and PU Respectively, “OOM” Denotes “Out of Memory”, and $\bullet$/$\circ$ Indicates whether SPGCC is Statistically Superior/Inferior to the Compared Methods According to the Pairwise $t$-Test at 0.05 Significance Level.

### v2

@@ Abstract
Hyperspectral images (HSI) clustering is an important but challenging task. The state-of-the-art (SOTA) methods usually rely on superpixels, however, they do not fully utilize the spatial and spectral information in HSI 3-D structure, and their optimization targets are not clustering-oriented. In this work, we first use 3-D and 2-D hybrid convolutional neural networks to extract the high-order spatial and spectral features of HSI through pre-training, and then design a superpixel graph contrastive clustering (SPGCC) model to learn discriminative superpixel representations. Reasonable augmented views are crucial for contrastive clustering, and conventional contrastive learning may hurt the cluster structure since different samples are pushed away in the embedding space even if they belong to the same class. In SPGCC, we design two semantic-invariant data augmentations for HSI superpixels: pixel sampling augmentation and model weight augmentation. Then sample-level alignment and clustering-center-level contrast are performed for better intra-class similarity and inter-class dissimilarity of superpixel embeddings. We perform clustering and network optimization alternatively. Experimental results on several HSI datasets verify the advantages of the proposed SPGCC compared to SOTA methods. Our code is available at https://github.com/jhqi/spgcc.
@@ I Introduction
Hyperspectral remote sensing techniques employ sensors to collect the reflectance of land-cover materials in hundreds of narrow and contiguous spectral bands, generating hyperspectral images (HSI), which have been widely applied in various fields, such as agricultural monitoring[1], mineral exploration[2] and urban planning[3]. Grouping HSI into several classes is essential in many applications[4, 5, 6, 7]. Recent state-of-the-art methods have achieved high accuracy in the supervised HSI classification task[8, 9, 10, 11]. However, due to the high cost of manual annotating, obtaining pixel-level labels is difficult.
Unsupervised clustering methods[12, 13, 14, 15] are gaining increasing attention. These methods aim to mine the relationships between pixels in the absence of labels and assign similar pixels to the same cluster. For example, center-based clustering methods, e.g., K-means[16] and fuzzy c-means[17], assume that instances have a spherical cluster structure, which may be not satisfied in the raw feature space of HSI. Subspace-based clustering methods, e.g., sparse subspace clustering[18], low-rank subspace clustering[19], and dense subspace clustering[20], divide instances into different low-dimensional subspaces, which are susceptible to redundant spectral features. Additionally, considering the large quantity of HSI pixels, subspace clustering methods may lead to unbearable computational costs because they need to calculate the self-represented coefficient matrix with time complexity of $\mathcal{O}(n^{3})$, where $n$ is the number of pixels. Therefore, some competitive subspace-based methods are limited to experimenting on a subset of the complete HSI[21, 22].
@@
To overcome the aforementioned drawbacks, we propose a superpixel-level clustering method for HSI with pixel-level pre-training, named superpixel graph contrastive clustering (SPGCC). First, we exploit the 3-D structure of HSI by dividing the original HSI into 3-D cubes. Then, 3-D and 2-D hybrid convolutions are performed on pixel cubes to extract the high-order spatial and spectral features of HSI simultaneously which can be preserved in superpixel-level representations. Next, we aggregate the topological information of superpixels from neighborhoods by GCN. To learn the discriminative and clustering-friendly superpixel representations, we utilize contrastive learning (CL), which brings samples close to their augmented views in the embedding space while pushing different samples’ representations far apart[32, 33, 34]. Proper data augmentations are crucial for CL, as augmented views and original samples should share the same semantics. To this end, we design two effective data augmentations for superpixels, i.e., pixel sampling augmentation and model weight augmentation, to obtain semantic-invariant augmented views. Additionally, traditional CL-based clustering methods treat different samples as negative pairs and push their representations to be far apart from each other[35], even for samples in the same class, which is harmful to clustering accuracy. Differently, we perform sample-level alignment and clustering-center-level contrast on superpixels to obtain better intra-class similarity and inter-class dissimilarity. To mitigate the impact of outliers, we select samples with high confidence to recompute clustering centers. In the whole framework, we alternatively conduct clustering and network optimization, using clustering results to guide the optimization. The proposed SPGCC has an improvement of 2.5%, 2.4%, and 2.7% over the best compared methods on three different HSI datasets.
@@
    We perform pixel-level pre-training with 3-D and 2-D hybrid CNN to extract high-order spatial and spectral information of HSI. Moreover, we separate it with the clustering network based on GCN, which enables our method to handle the large-scale HSI clustering task.
@@
    Two types of semantic-invariant data augmentations are proposed for HSI superpixels, which can improve contrastive learning by obtaining reliable positive samples.
@@
    Sample-level alignment and clustering-center-level contrast are combined to address the limitation of traditional contrastive learning methods that treat different samples of the same class as negative pairs, resulting in more clustering-friendly representations.
The rest of this paper is organized as follows. Section II briefly reviews preliminary about HSI pixel convolution, graph contrastive learning and HSI data augmentation. Section III presents the proposed SPGCC algorithm in detail, including pixel-level pre-training, graph construction, semantic-invariant superpixel augmentations, and graph contrastive clustering. In Section IV, we provide experimental results and analyses on several HSI datasets. Finally, we summarize our work in Section V.
@@ II-A HSI Pixel Convolution
where $d_{l-1}$ is the number of channels in the (l1)(l-1)-th layer, $\omega_{i,j,\tau}\in\mathbb{R}^{(2\delta+1)\times(2\gamma+1)\times({2\eta+1})}$ is the kernel’s weight parameter in the $j$-th channel of the $i$-th layer related to the $\tau$-th channel of the last layer, superscripts $(\sigma,\rho,\lambda)$ are indices along corresponding dimensions, $b_{i,j}$ is the bias, and $\phi$ is the activation function.
@@ II-B Graph Contrastive Learning
Several works have applied contrastive learning for graph data due to its remarkable representation ability[41, 42, 43]. Here, we focus on node-level graph contrastive learning (GCL) and divide it into three parts: graph data augmentation, graph encoder network, and contrastive loss function. A graph is denoted as $g=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ is the set of $N$ nodes and $\mathcal{E}$ is the set of $M$ edges. The graph structure is represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ and node attributes are denoted as $\mathbf{X}\in\mathbb{R}^{N\times d}$. We use $\Gamma$ to denote a set of different graph data augmentations and then augmented views are defined as:
@@
where $sim(\cdot)$ is the similarity function and $\tau$ is the temperature scaling parameter.
II-C Data Augmentation for HSI
Data augmentation can enhance the quality of original samples or generate additional reliable samples to improve training, which has been widely used in HSI tasks [45, 46, 40]. Here we mainly focus on the impact of data augmentation on superpixel-level graph contrastive learning.
For graph-structured superpixel data with node attributes and the adjacency matrix, general graph data augmentations can be directly applied, such as node mask, node noise, node shuffling, edge perturbation [47, 48]. However, these methods mainly rely on random perturbations of graph nodes or edges, which are uncontrollable and may erroneously alter semantic information of the graph. In contrastive learning, the augmented views are supposed to share the same semantic information with the original samples [33, 32], so the model can learn the common latent features by reducing the distance between original samples and corresponding augmented views in the representation space. Improper data augmentations may lead to negative representation learning. To this end, semantic-invariant data augmentations are crucial for conducting contrastive learning on HSI, which have been rarely investigated.
@@ III Proposed Method
Original HSI is 3-D structured and can be denoted as $\mathcal{S}\in\mathbb{R}^{H\times{W}\times{C}}$ with resolution of $H\times{W}$ and CC spectral bands. The target is to assign the total $N$ pixels into $K$ classes where $N=H\times{W}$. Our method consists of pixel-level pre-training and superpixel-level contrastive clustering. Since clustering is performed for superpixels, pixels’ labels are inherited from the belonged superpixels. The overall architecture of the proposed method is shown in Fig. 1. First, we employ a variational autoencoder composed of hybrid 3-D and 2-D CNNs for pixel-level pre-training, extracting high-order spatial and spectral features from HSI pixels. Subsequently, superpixel segmentation is performed and the graph structure of superpixels is constructed. In superpixel graph contrastive clustering, we propose two semantical-invariant data augmentations for superpixels and employ sample-level alignment together with clustering-center-level contrast as the target of network optimization. Clustering results are obtained by K-means and then high-confidence clustering centers are recomputed. We conduct clustering and network optimization alternatively.
@@
Encoder DecoderDecoder
Layer Kernel Output LayerLayer Kernel Output
Input - [-1,1,30,27,27] InputInput - [-1,128]
Conv3D [8,7,3,3] [-1,8,24,25,25] FCFC [128,256] [-1,256]
BN3D - - FCFC [256,23104] [-1,23104]
Conv3D [16,5,3,3] [-1,16,20,23,23] ReshapeReshape - [-1,64,19,19]
@@ III-C Superpixel Semantic-Invariant Data Augmentations
The second superpixel augmentation, model weight augmentation (MWA), is enlightened by the multi-head idea[51, 52]. We design dual branches in the final layer of the GCN encoder to obtain augmented views with semantic consistency. Specifically, the $L$-th layer contains two branches with unshared parameter matrices $\mathbf{W}_{(L)}^{1}$ and $\mathbf{W}_{(L)}^{2}$ which have the same architecture. Due to the different random initialization, two branches’ outputs $\mathbf{H}_{(L)}^{1}$ and $\mathbf{H}_{(L)}^{2}$ represent two different views of the input superpixel but with the same semantic, i.e., a further aggregation result based on the (L1)(L-1)-th layer. Considering the parameter efficiency and the diversity of positive samples, parameters are only unshared in the final layer, while shared in previous layers.
Combining PSA with MWA, we can obtain different augmented views of the superpixel. Specifically, representations of superpixels $\mathbf{X}^{sp}$ and the corresponding sampled pixels $\mathbf{X}^{p}_{sampled}$ are the inputs of the graph encoder. In the first L1L-1 layers, embeddings of superpixels and sampled pixels are calculated by Eq. (11), denoted as $\mathbf{H}_{(L-1)}^{sp}$ and $\mathbf{H}_{(L-1)}^{p}$ respectively. In the last layer, $\mathbf{H}_{(L-1)}^{sp}$ and $\mathbf{H}_{(L-1)}^{p}$ are inputted into both two branches, thus the outputs include four embeddings: $\mathbf{Z}^{sp_{1}},\mathbf{Z}^{sp_{2}},\mathbf{Z}^{p_{1}},\mathbf{Z}^{p_{2}}$ where $\mathbf{Z}^{sp_{1}}$ and $\mathbf{Z}^{p_{1}}$ are the outputs of the first branch for superpixels and sampled pixels, while $\mathbf{Z}^{sp_{2}}$ and $\mathbf{Z}^{p_{2}}$ are of the second. Subsequently, we normalize each row of these embeddings with $l_{2}$-norm.
@@
0: HSI $\mathcal{S}$; Superpixel segmentation $\mathbf{T}$; Max iteration II.
@@ III-D Superpixel Contrastive Clustering
Traditional contrastive clustering methods suffer from low intra-class consistency caused by wrong negative pairs[53]. To overcome this deficiency, we utilize sample-level alignment (SLA) and clustering-center-level contrast (CLC) as optimization objectives. Briefly, at the sample level, we do not treat different superpixels as negative pairs since they may belong to the same class. Superpixel embeddings are only encouraged to be aligned with their semantic-invariant augmented views. At the clustering center level, for a certain clustering center, it is natural to set the other K1K-1 centers as its negative views and perform contrast. Combining SLA with CLC, we hope to get superpixel embeddings with better intra-class similarity and inter-class dissimilarity.
@@
$\mathcal{L}_{CLC}=\frac{1}{K}\sum_{k=1}^{K}-log\frac{exp(\mathbf{c}_{k}^{1}\cdot\mathbf{c}_{k}^{2}/\tau)}{\sum_{j=1}^{K}exp(\mathbf{c}_{k}^{1}\cdot\mathbf{c}_{j}^{1}/\tau)},$ (19)
where $\tau$ is the temperature scaling parameter. In practice, we adopt the symmetric form of Eq. (19) considering the equal importance of the first and the second views.
@@
III-E Computational Complexity Analysis
There are several learning stages in SPGCC. At pixel-level pre-training, it takes $\mathcal{O}(\delta^{3}\gamma^{3})$ and $\mathcal{O}(\delta^{2}\gamma^{2})$ for each layer of 3-D and 2-D CNNs, where $\delta$ and $\gamma$ denote the length of feature maps and kernels respectively. It takes $\mathcal{O}(NlogN)$ to perform ERS segmentation on HSI, where $N$ is the number of pixels. For GCN, let $\|\mathbf{A}\|_{0}$ be the number of non-zero elements in the adjacency matrix, $L$ be the number of layers, $d$ be the node feature dimension and assumed to be fixed. It takes $\mathcal{O}(L\|\mathbf{A}\|_{0}d+LMd^{2})$ to aggregate superpixels in the neighbourhood, where $M$ is the number of superpixels. Obtaining clustering results with K-means takes $\mathcal{O}(TKMd)$, where $K$ and $T$ are the number of classes and iterations.
@@ IV-A Datasets
Dataset Size #Band #Class #Labeled pixels
India Pines (145, 145) 200 16 10249
Salinas (512, 217) 204 16 54129
Pavia University (610, 340) 103 9 42776
@@ IV-B Experiment Setup
1) Implementation Details: Our method consists of pixel-level pre-training and superpixel-level contrastive clustering. For the first stage, we divide 3-D pixel cubes with a fixed window size of 27 $\times$ 27 and reduce the spectral dimension to 30 on India Pines, 15 on Salinas, and 15 on Pavia University for computational efficiency. The architecture of CNNs is the same as shown in Table I for three datasets. Adam optimizer is utilized. We set the learning rate and weight decay to $1\text{\times}{10}^{-3}$ and $5\text{\times}{10}^{-4}$ respectively. When the pre-training of the VAE module is finished, all the pixel cubes are mapped to feature vectors and saved for the following learning stage. For the second stage, we use GCN as the graph encoder, where the sizes of the hidden layer and the output layer are 1024 and 512 respectively. We perform K-means on superpixel embeddings to obtain clustering assignments per five epochs. The temperature scaling parameter of contrast is fixed to 0.5. The weight of clustering-center-level contrast loss is fixed to 0.1. We tune the learning rate $\eta$ from $1\text{\times}{10}^{-5}$ to $1\text{\times}{10}^{-1}$. Other hyperparameters are shown in Table III. All the experiments are repeated 10 times and we report the average value and standard deviation of clustering metrics finally.
@@
Dataset $M$ $\eta$ $\alpha$ $\lambda$ $L$
India Pines 1100 $1\text{\times}{10}^{-5}$ 0.1 0.75 3
Salinas 2700 $1\text{\times}{10}^{-5}$ 0.1 0.55 3
Pavia University 2200 $1\text{\times}{10}^{-4}$ 0.1 0.25 3
2) Compared Methods: To comprehensively evaluate the performance and applicability of the proposed SPGCC, we introduce ten baselines for comparison, including K-means[16], fuzzy c-means (FCM)[17], sparse subspace clustering (SSC)[18], scalable sparse subspace clustering (SSSC)[54], selective sampling-based scalable sparse subspace clustering (S5C)[55], superpixelwise PCA approach (SuperPCA)[56], graph convolutional optimal transport (GCOT)[57], neighborhood contrastive subspace clustering (NCSC)[13], superpixel-level global and local similarity graph-based clustering (SGLSC)[12], dual graph autoencoder (DGAE)[14], spectral–spatial and superpixelwise unsupervised linear discriminant analysis (S3-ULDA)[58]. Specifically, for K-means and FCM, we perform clustering at the pixel level. For SSC, SSSC, S5C, clustering is performed at the superpixel level to overcome the high memory cost of the subspace clustering. For other methods, we follow the original settings proposed in their papers. A brief summarization of the evaluated methods can be found in Table IV.
@@
Method Superpixel Subspace clustering Graph learning Deep learning Dimension reduction
@@
SSC $\checkmark$ $\checkmark$
SuperPCA $\checkmark$ $\checkmark$
SSSC $\checkmark$ $\checkmark$
S5C $\checkmark$ $\checkmark$
@@
SGLSC $\checkmark$ $\checkmark$ $\checkmark$
DGAE $\checkmark$ $\checkmark$ $\checkmark$
NCSC $\checkmark$ $\checkmark$ $\checkmark$
S3-ULDA $\checkmark$ $\checkmark$
SPGCC $\checkmark$ $\checkmark$ $\checkmark$
@@
3) Clustering Metrics: To quantitatively evaluate the clustering performance, nine popular metrics are utilized: overall accuracy (OA), average accuracy (AA), kappa coefficient (Kappa), normalized mutual information (NMI), adjusted Rand index (ARI), F1-score (F1), Precision, Recall and Purity. Among these metrics, OA, AA, NMI, F1, Precision, Recall, and Purity range in [0%,100%] while Kappa and ARI range in [-100%, 100%]. For each metric, a higher score indicates a better clustering result. Similar to the common process of clustering evaluation, the Hungarian algorithm [59] is adopted to map the predicted labels according to the ground truth before calculating metrics. In addition, we log the time cost of each method.
@@
TABLE V: Clustering Performance on HSI Datasets, India Pines, Salinas and Pavia University are Abbreviated to IP, SA and PU Respectively, “OOM” Denotes “Out of Memory”, and $\bullet$$\circ$ Indicates whether SPGCC is Statistically Superior/Inferior to the Compared Methods According to the Pairwise $t$-Test at 0.05 Significance Level.